# X-ray2CTPA: Generating 3D CTPA scans from 2D X-ray conditioning

### v3

@@ Discussion
The purpose of this work was to generate 3D CTPA scans from paired 2D CXRs. To address this problem, we used a novel 3D diffusion model with BiomedCLIP Visual encoder, additional adversarial guidance, classifier and perceptual loss. We had many challenges when implementing this project: the number of samples in our dataset is relatively small, the size of each CTPA scan is large, the 3D nature of the CT scans enforces consistency throughout the generated 3D scan and between 2D slices which makes it hard to use the common 2D pretrained models. In addition, the 2D CXR data inputs are not registered or aligned to the CTPA scans. Finally, the application of PE is challenging due to the small size of the clots in the scans.
@@
We were able to generate somewhat realistic, consistent and anatomically correct 3D scans. The generated 3D scans are not identical to the ground truth CTPAs but are similar and capture their structure. We experimented in the image size and additional modifications to the model and loss functions. Our best performing model improved the CXR PE classification results by 11%; Improving from an AUC of 0.691 [95% CI: 0.54-0.77] to 0.803 [95% CI: 0.72-0.9]. In addition, the specificity of the model improved, which in a clinical setting may potentially lead to a reduction of redundant CTPA scans performed for PE suspected patients. Unlike other diffusion studies in the medical field, this research uniquely applies generated data to address a real-life clinical challenge rather than focusing on the quality of the generated samples. As far as we know, this is the first model that attempts to convert 2D chest X-rays to CTPA scans. We also demonstrate the model’s generalizability by fine-tuning it on an extremely small number of samples to achieve 2D CXR to 3D CT translation.
@@
This study is subject to a few significant limitations. It was designed as a single-center, retrospective study, which is associated with well-known shortcomings and inherent limitations. In addition, generalizing the methods and results of our study to other predictive tasks with different modalities should be examined carefully, as our comparison and analysis were limited to CXR and CTPA modalities and to the task of PE classification. Finally, while our model generates 3D CTPA scans which are close to the ground truth, they are by no means identical to them and currently cannot be used to replace CTPA scans. This is shown throughout the paper, specifically, in the Failed examples Section, where we present examples of failure and model hallucinations. We hypothesize that in order for the generated scans to be used as actual replacements for the original CTPAs, we need a much larger dataset.
@@ Methods
We used the pretrained version of a 2D-VAE that was fine-tuned from the original kl-f8 autoencoder to improve the reconstruction of faces [30, 31]. The model remained frozen without any further fine-tuning for our dataset. The 3D-CTPA scans were fed to the 2D-VAE sequentially slice by slice and then concatenated before being fed into the 3D diffusion model. For the conditional setting, we added 2D CXR scans as conditioning to the diffusion model. We hypothesize that the CXR to CTPA modality shift is more like text to image conversion rather than image-to-image inpainting. Therefore, we applied the conditioning by concatenation of the CXR embedding throughout the 3D U-net layers, rather than applying it as concatenation to the noise as input of the U-net. The CXR was encoded to a 512-vector using BiomedCLIP Visual encoder [32]. We further improve the 3D scan generation by imposing a 4 layered 3D discriminator at the output of the diffusion process. In addition, a DenseNet3D-121 [33] classification was also applied on the generated latents to introduce label information and improve classification results.
The 2D-VAE and BiomedCLIP Visual encoder remained frozen during the training and only the 3D diffusion model and the 3D discriminator were trained. Due to the small number of samples, we pretrained the diffusion model on the largest publicly available annotated PE dataset (RSPECT) - 2020 RSNA Pulmonary Embolism Detection Kaggle Challenge, which is comprised of more than 7,000 3D CTPA studies from five international research centers [18].
@@
Figure 7: X-ray2CTPA model architecture: Black arrows - Training: The CTPA scan are fed slice by slice to the 2D-VAE in a sequential manner and compressed by a factor of 8. The compressed slices are concatenated. The latent representation is noised in the forward diffusion process and is then used as input to the 3D denoising U-net model. The CXR acts as conditioning to the model, fed into the CLIP Vision encoder. Bottom - Sampling: A noise latent and CXR are the inputs to the model and the output is the reconstructed 3D CTPA scan.
@@
Our cohort was constructed from 898 patients who underwent both a chest X-ray scan and a chest CT scan with contrast agent - CT pulmonary angiography (CTPA) within a 24-hour difference and from a single center. The chest CT volumes were annotated by board certified radiologists for a single label determining whether the patients had pulmonary embolism (PE) or not. The CTPA scans were anonymized and categorized by expert opinion indicating whether PE existed or not (no severity indication and no marked PE segmentation) according to the radiologist’s report. For the evaluation of the models, we used a stratified sixfold cross-validation. 305 of the samples (34 %) were marked positive for PE. The dataset specifications are presented in Table 5. An institutional review board (IRB) approval was granted for this work. Informed consent was waived by the IRB committee.
@@
Train $803$ ($89$)
Validation $50$ ($6$)
Test $45$ ($5$)
@@
The 3D CTPA scans had a spacial resolution of 512×512 with varying number of axial slices. All scans were resampled to a standard uniform voxel spacing of 1 mm in all dimensions. The pixel values were converted to Hounsfield units (HU) and windowed to the range of $(-100HU,+900HU)$, to enhance the lower and upper boundaries of the area of interest limits of the HU scale. Finally, the scans were cropped to the region of the lung area using lung segmentation and resized to a size of 256x256x128 voxels (height x width x depth). Due to this preprocessing step, the axial slices of the scans seem a bit deformed. We min–max normalized the scans to the range of ($-1$,$1$). To achieve a reduced length without additional spacial manipulation of the data, the scans were divided into two segments of 256x256x64 each by selecting every alternate slice. Additionally, we augmented all datasets by vertically flipping the images during training with a probability of $50\%$.
The 2D CXR scans were normalized and resized to the size of 224 x 224 x 3 to match the size of the BiomedCLIP Vision encoder [32].
@@
We utilize multiple loss functions to improve its training, each with a distinct purpose. The overall loss is a combination of $l_{1}$, perceptual similarity (LPIPS) [24] weighted by $\lambda_{1}$, a discriminator loss, weighted by $\lambda_{2}$ and a classification loss weighted by $\lambda_{3}$. $l1$ loss ensures the generated images closely match the ground truth. LPIPS, leveraging the 2D ResNet-50 based RadImageNet architecture [27], enhances the model’s grasp of complex semantic details and ensures perceptual similarity. The discriminator loss further ensures the generated images’ realism, and the classifier loss guides the samples to the PE classification task. $\lambda_{1},\lambda_{2},\lambda_{3}$ were set to $0.01,0.001$ and $0.08$ respectively. The loss can be summarized as follows:
@@
The performance evaluation on the test set covered AUC, sensitivity (also known as recall), specificity, accuracy, the positive predictive value (PPV, also known as precision), and the negative predictive value (NPV). The DeLong method [26] was used to calculate the 95% confidence intervals for the AUC. The probability thresholds for predicting positive samples were determined by the sensitivities and specificities on the validation set, which ensured high specificities while keeping reasonable sensitivities. These thresholds were determined by Youden’s index [34], which finds the optimal joint sensitivity and specificity.
@@
All models were trained on a single NVIDIA RTX A5000 GPU, and each took about a week of training. The code for our model was based on Medical Diffusion [11] with various alterations. The training parameters included a base learning rate of $1e^{-4}$ for the unconditional setting and when pretraining using RSPECT dataset and a learning rate of $1e^{-5}$ when fine-tuned for the final conditional model. An additional pretraining step was added in which we used the RSPECT dataset for pretraining but in a conditional setting. We used artificial CXR images or digitally reconstructed Radiographs (DRRs) as conditioning. The DRRs were created using DiffDRR [35] Python package for differentiable X-ray rendering. All were trained with a batch size of 2 with AdamW optimizer and Cosine Anneal scheduler. We used DDPM sampling with 1000 steps.
Acknowledgements
The research in this publication was supported by the Israel Science Foundation (ISF) under grant number 20/2629.
@@ Acknowledgements
Funding. This study was funded by the Israel Science Foundation (ISF) under Grant number 20/2629.
@@
Consent for publication. The IRB committee of Sheba Medical Center waived the need of informed consent.
Data availability. The CXR-CTPA dataset generated and/or analysed during the current study are not publicly available due patient privacy regulations but are available from the corresponding author on reasonable request. The LIDC-IDRI (https://wiki.cancerimagingarchive.net/pages/viewpage.action?pageId=1966254) dataset is publicly available at the cancer imaging archive (TCIA) [36].
Author contributions statement
Guarantor of integrity of the entire study, planning and supervision, E.K, G.A, H.G.; data curation or data analysis/interpretation, E.K, Y.B.; study concepts/study design and methodology, N.C, R.G, H.G.; software, visualization and writing of the original draft N.C.; Clinical consulting G.A., E.K.; manuscript drafting or manuscript revision for important intellectual content, all authors.; approval of final version of submitted manuscript, all authors.; agrees to ensure any questions related to the work are appropriately resolved, all authors.; literature research, N.C., Y.B.; experimental studies, N.C, R.G, H.G.; statistical analysis, N.C., R.G., H.G.; and manuscript editing, R.G, N.C, E.K, E.K, Y.B, H.G.

### v4

@@ Discussion
The purpose of this work was to generate 3D CTPA scans from paired 2D CXRs in order to enhance PE classification from CXR images. To address this problem, we used a novel 3D diffusion model with BiomedCLIP Visual encoder, additional adversarial guidance, classifier and perceptual loss. We had many challenges when implementing this project: the number of samples in our dataset is relatively small, the size of each CTPA scan is large, the 3D nature of the CT scans enforces consistency throughout the generated 3D scan and between 2D slices which makes it hard to use the common 2D pretrained models. In addition, the 2D CXR data inputs are not registered or aligned to the CTPA scans. Finally, the application of PE is challenging due to the small size of the clots in the scans.
@@
Our best performing model improved the CXR PE classification results by 11%; Improving from an AUC of 0.691 [95% CI: 0.54-0.77] to 0.803 [95% CI: 0.72-0.9]. In addition, the specificity of the model improved, which in a clinical setting may potentially lead to a reduction of redundant CTPA scans performed for PE suspected patients. Unlike other diffusion studies in the medical field, this research uniquely applies generated data to address a real-life clinical challenge rather than focusing on the quality of the generated samples. We were able to generate somewhat realistic, consistent and anatomically correct 3D scans. The generated 3D scans are not identical to the ground truth CTPAs but are similar and capture their structure. We experimented with the image size and additional modifications to the model and loss functions. As far as we know, this is the first model that attempts to convert 2D chest X-rays to CTPA scans. We also demonstrate the model’s generalizability by fine-tuning it on an extremely small number of samples to achieve 2D CXR to 3D CT translation.
The classification results presented not only substantiates the findings of this study but also highlights the value of the proposed methodology. We present Grad-CAM visualizations that demonstrate our classifier focuses on anatomically and clinically relevant regions indicative of pulmonary embolism rather than on potential hallucinated features. Furthermore, our initial human observer study shows that the availability of generated 3D CTPA scans significantly enhances PE detection from CXR, underscoring the potential of these advanced imaging methods to improve diagnostic confidence when standard CXR imaging is insufficient. We argue that the generation task acts as a valuable pre-text task for classification, serving as a form of regularization rather than the primary output. This perspective aligns with OpenAI’s Image-GPT [33], which demonstrates that generative image modeling is a promising route to learn high-quality unsupervised image representations. Furthermore, performing the classification within the 3D diffusion latent space reinforces this assertion.
We acknowledge the growing number of studies in the field of medical generative AI, particularly in Chest X-ray to CT translation. However, our work distinguishes itself in several key aspects: 1. To the best of our knowledge, this is the first study to address the challenge of converting CXRs to CT scans directly, rather than relying on DRRs. Unlike DRRs, which contain significantly more information and are inherently aligned with CT scans, CXRs present a greater challenge due to their limited information and lack of alignment. Our proposed method addresses this challenge without requiring extensive preprocessing or alignment between modalities, resulting in a more robust and adaptable approach. 2. Many generative models in this field are primarily focused on data generation and are rarely applied to solving specific downstream tasks, such as classification or segmentation. In contrast, this paper goes beyond data augmentation by directly demonstrating the utility of our approach in solving a classification problem, showcasing its practical application and impact.
We have tested multiple CXR-focused baselines, and were not able to compete with the generative-based results presented. On the training set these models reach a performance comparable to our pipeline; its shortfall appears once we test on unseen and the model’s ability to generalize. The key difference is geometric context: a single CXR provides only a 2D projection, whereas the diffusion step generates a 3D representation of the lungs. This volumetric view allows the downstream classifier to inspect clot-relevant patterns that are entirely collapsed in a projection image, leading to better generalization. Similar trends appear in other fields, in which introducing a generative step can provide the model with richer or more varied information than the raw inputs alone.Some examples are: ImageGPT [33], shows that diffusion-style pretext tasks learn high-quality unsupervised latents that boost downstream accuracy; GAN-based synthetic augmentation in medical imaging routinely improves classification performance [34, 35]. Likewise, in NLP, generative pre-training of language models such as BERT and GPT substantially lifts text-classification benchmarks [36]. These generative steps expand the discriminative feature space, enhance generalization, and ultimately yield representations that surpass direct classification on raw images.
@@
This study is subject to a few significant limitations. It was designed as a single-center, retrospective study, which is associated with well-known shortcomings and inherent limitations. In addition, generalizing the methods and results of our study to other predictive tasks with different modalities should be examined carefully, as our comparison and analysis were limited to CXR and CTPA modalities and to the task of PE classification. Finally, while our model generates 3D CTPA scans which are close to the ground truth, they are by no means identical to them and currently cannot be used to replace CTPA scans.
While X-ray2CTPA model can generate high-quality 3D CTPA samples that closely resemble the ground truth in structure and appearance, there are instances where it slightly diverges or fails to accurately replicate the ground truth. Some examples from our test set are presented in Fig. 9. The main areas where the model fails are around small veins or when rare abnormalities or additional pathologies (other than PE) are presented. In some of these cases the model generates the pathology in some variant but may replicate it in a wrong location or fail to generate the pathology whatsoever. We hypothesis that this issue may be caused by the high variability nature of the CTPA samples and due to the fact, that in many cases PE is secondary to other patient comorbidities. We believe that training the model with a larger dataset may solve this issue in future work. We hypothesize that in order for the generated scans to be used as actual replacements for the original CTPAs, we need a much larger dataset.
We acknowledge that, at its current performance level, our model does not meet the stringent sensitivity standards typically required for clinical triage. Rather, our work represents an initial proof of concept demonstrating the potential utility of generated 3D CTPA scans in guiding preliminary decisions. For context, the widely used Wells Score achieves an AUC of approximately 0.78 [37], so our X-ray2CTPA approach could serve as an adjunct - augmenting the Wells Score rather than replacing CT imaging - to refine pre-test probability and guide the decision to perform a D-dimer assay from a simple chest X-ray. As the technology matures, incorporating larger datasets, improved architectures, and refined training strategies may increase both sensitivity and specificity. This incremental progress could eventually enable the model to serve as a practical triage support tool, flagging likely PE cases for more urgent follow-up imaging while reducing unnecessary scans in low-risk scenarios Fig. 10 illustrates a proposed workflow combining Wells Score, generated CTPA, and D dimer testing.
Figure 10: A possible clinical workflow in which X-ray2CTPA can be used. The X-ray2CTPA model can be used along with the Wells Score for decision whether to perform D-Dimer or not. In this setup the model serves as a practical triage support tool (Icon images from Microsoft PowerPoint 2023.).
@@ Methods
The 3D-CTPA scans were fed to the 2D-VAE sequentially slice by slice and then concatenated before being fed into the 3D diffusion model. For the conditional setting, we added 2D CXR scans as conditioning to the diffusion model. We hypothesize that the CXR to CTPA modality shift is more like text to image conversion rather than image-to-image in-painting. Therefore, we applied the conditioning by concatenation of the CXR embedding throughout the 3D U-net layers, rather than applying it as concatenation to the noise as input of the U-net. The CXR was encoded to a 512-vector using BiomedCLIP Visual encoder [38]. We further improve the 3D scan generation by imposing a 4 layered 3D discriminator at the output of the diffusion process. In addition, a DenseNet3D-121 [39] classification was also applied on the generated latents to introduce label information and improve classification results.
@@
Figure 11: X-ray2CTPA model architecture: Black arrows - Training: The CTPA scan are fed slice by slice to the 2D-VAE in a sequential manner and compressed by a factor of 8. The compressed slices are concatenated. The latent representation is noised in the forward diffusion process and is then used as input to the 3D de-noising U-net model. The CXR acts as conditioning to the model, fed into the CLIP Vision encoder. Bottom - Sampling: A noise latent and CXR are the inputs to the model and the output is the reconstructed 3D CTPA scan.
PE Classification.
We employed two single-modality encoders for PE classification. One encoder processes CTPA data, while the other handles CXR data. These encoders serve as baseline classification models. In addition, we leverage the trained CTPA encoder as a pretrained model to classify PE on the CTPA scans generated by our X-ray2CTPA model.
CTPA Classifier.
The trained model is introduced during the training phase of the X-ray2CTPA model as a classification loss, and it is subsequently employed to evaluate the generated CTPA scans once training is complete. In addition, the classifier is applied to the generated latents during training to inject label information and enhance CXR classification results.
We encountered several challenges that guided our selection of the CTPA classifier architecture. Our dataset is relatively small and provides only a single binary label for each entire 3D volume. We were further constrained by the diffusion model’s ability to generate only a maximum of 64 frames along the z-axis. We explore the optimal CTPA classifier for both real and generated scans. We applied two complementary approaches, one operating directly in pixel space and the other in the diffusion model’s latent space: 1. INSPECT pixel-space classifier [30]: Although INSPECT’s published performance is not state-of-the-art (SOTA) for PE classification, it embodies the current best-practice strategy for this task. The model is pre-trained with 2D label supervision for every slice in a 3D volume, and can process an unlimited number of slices without resampling. The model is comprised of a two-phase pipeline: At first, a 2D ResNet-101 is trained using 2D labels from each slice of the 3D scan. This trained model is then frozen and used as a feature extractor. In the second phase, the extracted features are fed into a sequential LSTM model for a final prediction. We note that we were unable to fine-tune the ResNet backbone on our dataset, as our labels are provided only at the 3D scan level. As a result, we could only train the LSTM part of the model. 2. Latent space classifier: We also trained a 3D DenseNet-121 on the latent representations produced by the 2D VAE that is used in our diffusion model. This network operates on the compressed (scaled-down) version of each CTPA volume, enabling efficient learning while preserving volumetric context.
For the latent classifier, in the first stage, the classifier was trained on the original CTPA dataset (excluding our held-out test set) using a scaled-down latent representation derived through the same processing pipeline as the generated latents from the X-ray2CTPA diffusion model. Once the classifier was trained, we incorporated it into the X-ray2CTPA model as a classification loss. During inference, we generated CTPA latents for the entire dataset and fine-tuned the same model for evaluation.
CXR Classifier.
We evaluated several pre-trained models for the CXR encoder and fine-tuned them for the specific task of PE classification: 1. RadImageNet [40], a ResNet-50 based model pretrained on millions of radiologic images. 2. BiomedCLIP Vision encoder [38]. The BiomedCLIP visual encoder, part of the BiomedCLIP foundation model, was pretrained on a biomedical dataset of 15 million figure-caption pairs from PubMed Central. It achieves SOTA performance across a variety of medical vision-language processing (VLP) tasks, including cross-modal retrieval, image classification, and visual question answering. This is the same encoder that was used in the X-ray2CTPA model for CXR conditioning. 3. DINOv2 [41], one of the most powerful existing 2D foundation model. DINOv2 is a self-supervised distillation method applied on Vision Transformers (ViTs)  [42]. The model’s self-supervised ViT features contain explicit information about the semantic segmentation of an image and the extracted features can be used for many downstream tasks. The best performing encoder was the RadImageNet model, therefore we chose it as our baseline.
Training Strategy
The diffusion model was pretrained on the RSPECT dataset [43]. Then, the diffusion model was fine-tuned on our dataset in a two-stage setting. First, the model was fine-tuned in an unconditional manner, without any CXR conditioning. In the second stage, the CXR conditioning was added, and the model was further trained. Even though the datasets were relatively small, we found that each model converged and generated realistic synthetic images without a lot of hyperparameter finetuning. During training, the 2D-VAE and BiomedCLIP visual encoder remained frozen, while the 3D diffusion model, classifier, and discriminator were fine-tuned and trained further.
We utilized a pretrained 2D-VAE, fine-tuned from the original KL-f8 autoencoder, which demonstrated improved face reconstruction capabilities [44, 45]. Given its exceptional feature extraction capabilities on our dataset, the 2D-VAE required no additional fine-tuning and remained frozen throughout the training process. The BiomedCLIP visual encoder, part of the BiomedCLIP foundation model, was pretrained on a biomedical dataset of 15 million figure-caption pairs from PubMed Central. We selected BiomedCLIP for its SOTA performance across a variety of medical vision-language processing (VLP) tasks, including cross-modal retrieval, image classification, and visual question answering. For the discriminator, we employed the medical-pretrained version of MONAI Generative [46], using its Patch-GAN discriminator architecture [47], which was further fine-tuned during training.
To address the challenge of limited sample size, we pretrained the diffusion model on the RSPECT dataset - the largest publicly available annotated PE dataset. RSPECT, from the 2020 RSNA Pulmonary Embolism Detection Kaggle Challenge, comprises over 7,000 3D CTPA studies from five international research centers [43]. The classifier used in the X-ray2CTPA diffusion model pipeline was performed in the latent space and pretrained on latents extracted from the RSPECT dataset using the 2D-VAE. This classifier was trained separately before integration into the full X-ray2CTPA diffusion model setting.
@@
Our cohort was constructed from 898 patients who underwent both a chest X-ray scan and a chest CT scan with contrast agent - CTPA within a 24-hour difference and from a single center. The chest CT volumes were annotated by board certified radiologists for a single label determining whether the patients had PE or not. The CTPA scans were anonymized and categorized by expert opinion indicating whether PE existed or not (no severity indication and no marked PE segmentation) according to the radiologist’s report. During the anonymization process, each patient is assigned a unique ID that links their various tests and prevents data leakage. For the evaluation of the models, we used a stratified sixfold cross-validation. 305 of the samples (34 %) were marked positive for PE. The dataset specifications are presented in Table 8.
The institutional review board (IRB) of Sheba Medical Center approval was granted for this work. The need for informed consent was waived by the ethics committee. All methods and experiments were performed in accordance with relevant guidelines and regulations. The research was performed in accordance with the Declaration of Helsinki.
@@
Train $724$ ($80.6$)
Validation $88$ ($9.8$)
Test $86$ ($9.6$)
@@
The 3D CTPA scans had a spacial resolution of 512 × 512 with varying number of axial slices. All scans were resampled to a standard uniform voxel spacing of 1 mm in all dimensions. The pixel values were converted to Hounsfield units (HU) and windowed to the range of $(-100HU,+900HU)$, to enhance the lower and upper boundaries of the area of interest limits of the HU scale. Finally, the scans were cropped to the region of the lung area using lung segmentation and resized to a size of 256 × 256 × 128 voxels (height × width × depth). Due to this preprocessing step, the axial slices of the scans seem a bit deformed. We min–max normalized the scans to the range of ($-1$,$1$). To achieve a reduced length without additional spacial manipulation of the data, the scans were divided into two segments of 256 × 256 × 64 each by selecting every alternate slice. Additionally, we augmented all datasets by vertically flipping the images during training with a probability of $50\%$.
The 2D CXR scans were normalized and resized to the size of 224 × 224 × 3 to match the size of the BiomedCLIP Vision encoder [38]. Although the 2D CXR and 3D CTPA scans are paired, we did not use any pre-registration procedures in our pipeline.
@@
We utilize multiple loss functions to improve its training, each with a distinct purpose. The overall loss is a combination of $l_{1}$, LPIPS [24] weighted by $\lambda_{1}$, a discriminator loss, weighted by $\lambda_{2}$ and a classification loss weighted by $\lambda_{3}$. $l1$ loss ensures the generated images closely match the ground truth. LPIPS, leveraging the 2D ResNet-50 based RadImageNet architecture [40], enhances the model’s grasp of complex semantic details and ensures perceptual similarity. The discriminator loss further ensures the generated images’ realism, and the classifier loss guides the samples to the PE classification task. $\lambda_{1},\lambda_{2},\lambda_{3}$ were set to $0.01,0.001$ and $0.08$ respectively. The loss can be summarized as follows:
@@
The performance evaluation of the test set covered AUC, sensitivity (also known as recall), specificity, accuracy, the positive predictive value (PPV, also known as precision), and the negative predictive value (NPV). The DeLong method [18] was used to calculate the 95% confidence intervals for the AUC. The probability thresholds for predicting positive samples were determined by the sensitivities and specificities of the validation set, which ensured high specificities while keeping reasonable sensitivities. These thresholds were determined by Youden’s index [48], which finds the optimal joint sensitivity and specificity.
@@
All models were trained on a single NVIDIA RTX A5000 GPU, and each took about a week of training. The code for our model was based on Medical Diffusion [11] with various alterations. The training parameters included a base learning rate of $1e^{-4}$ for the unconditional setting and when pretraining using RSPECT dataset and a learning rate of $1e^{-5}$ when fine-tuned for the final conditional model. An additional pretraining step was added in which we used the RSPECT dataset for pretraining but in a conditional setting. We used artificial CXR images or DRRs as conditioning. The DRRs were created using DiffDRR [49] Python package for differentiable X-ray rendering. All were trained with a batch size of 2 with AdamW optimizer and Cosine Anneal scheduler. We used DDPM sampling with 1000 steps.
@@
Data availability. The CXR-CTPA dataset generated and/or analyzed during the current study are not publicly available due patient privacy regulations but are available from the corresponding author (Noa Cahan - E-mail: noa.cahan@gmail.com ) on reasonable request. The LIDC-IDRI (https://wiki.cancerimagingarchive.net/pages/viewpage.action?pageId=1966254) dataset is publicly available at the cancer imaging archive (TCIA) [50].
Acknowledgments. We thank Prof. Mayte Suarez-Farinas for her valuable support and guidance during this study. This research was supported by the Israel Science Foundation (ISF), grant no. 20/2629.
Author contributions statement. Guarantor of integrity of the entire study, planning and supervision, E.K, G.A, H.G.; data curation or data analysis/interpretation, E.K, Y.B.; study concepts/study design and methodology, N.C, R.G, H.G.; software, visualization and writing of the original draft N.C.; Clinical consulting G.A., E.K.; manuscript drafting or manuscript revision for important intellectual content, all authors.; approval of final version of submitted manuscript, all authors.; agrees to ensure any questions related to the work are appropriately resolved, all authors.; literature research, N.C., Y.B.; experimental studies, N.C, R.G, H.G.; statistical analysis, N.C., R.G., H.G.; and manuscript editing, R.G, N.C, E.K, E.K, Y.B, H.G.